OpenStudy (anonymous):

The velocity of an object moving along a straight line is v(t) = t^2-10 t+16. Find the displacement over the time interval [1, 7]. Find the total distance traveled by the object.

myininaya (myininaya):

her is a clue how do you get from velocity to position?

OpenStudy (anonymous):

integrate it

OpenStudy (anonymous):

1/3t^3 - 5t^2 + 16t

myininaya (myininaya):

yes

OpenStudy (anonymous):

so should i just put s(7) - s (1)

OpenStudy (anonymous):

that will give me displacement or distance ?

myininaya (myininaya):

That would give you displacement.

myininaya (myininaya):

you can think of displacement as like net distance

OpenStudy (anonymous):

ok how to find total distance ?

myininaya (myininaya):

the same way you would find area instead of net area

myininaya (myininaya):

\[v(t)=t^2-10t+16\] When is v positive and when is v negative on the interval 1 to 7?

OpenStudy (anonymous):

2 to 5 ?

myininaya (myininaya):

if you factor v we get (x-2)(x-8) ------|------|----- 2 8 positive negative positive So you will have to split your integral into two integrals one going from 1 to 2 and the other going from 2 to 7

myininaya (myininaya):

\[|\int\limits_{1}^{2}v dt | +| \int\limits_{2}^{7} v dt |\]

OpenStudy (anonymous):

ok so i integrate vt and get st then do s(2) -s(1) + s(7) - s(2) ?

myininaya (myininaya):

|s(2)-s(1)|+|s(7)-s(2)| <--correction

OpenStudy (anonymous):

oh ok thanks

myininaya (myininaya):

You might think this is useful: http://saylor.org/site/wp-content/uploads/2011/04/DISPLACEMENT-vs.pdf

OpenStudy (anonymous):

is displacement -ve ?

OpenStudy (anonymous):

i got -30 for displacement

myininaya (myininaya):

that's good

OpenStudy (anonymous):

but how can displacement be -ve ?

myininaya (myininaya):

I'm pretty that signifies the direction

myininaya (myininaya):

the negative part

myininaya (myininaya):

you are moving backwards

OpenStudy (anonymous):

and the total distance i got is -7.27

OpenStudy (anonymous):

am i doing something wrong

myininaya (myininaya):

yeah distance shouldn't be negative

myininaya (myininaya):

did you do |s(2)-s(1)|+|s(7)-s(2)| ?

OpenStudy (anonymous):

yes

myininaya (myininaya):

you do no it is impossible for absolute value of something to be negative right?

myininaya (myininaya):

know*

OpenStudy (anonymous):

yes so its 7.273

myininaya (myininaya):

so let me see if i can help you with the arithmetic

OpenStudy (anonymous):

that website say tha we should subtract the - part from +.....and we are adding both integrals

myininaya (myininaya):

\[s(t)=\frac{t^3}{3}-10 \frac{t^2}{2} +16t+C\] \[s(t)=\frac{t^3}{3}-5t^2+16t+C\] \[|s(2)-s(1)|=|(\frac{2^3}{3}-5(2)^2+16(2))-(\frac{1^3}{3}-5(1)^2+16(1)|\] \[|s(7)-s(2)|=|(\frac{7^3}{3}-5(7)^2+16(7))-(\frac{2^3}{3}-5(2)^2+16(2))|\] After you computer these Add them

myininaya (myininaya):

compute

myininaya (myininaya):

the absolute value around those differences mean you need to take the absolute value

myininaya (myininaya):

you are suppose to be adding a positive to a positive

OpenStudy (anonymous):

we have + and - so should it be s(+) - s(-)

myininaya (myininaya):

\[|\frac{8}{3}-20+32-\frac{1}{3}+5-16|+|\frac{343}{3}-245+112-\frac{8}{3}+20-32|\]

OpenStudy (anonymous):

oh ok so answer is 40.65 or 41 ?

myininaya (myininaya):

\[|\frac{7}{3}+12-11|+|\frac{335}{3}-133-12|\] \[|\frac{7}{3}+1|+|\frac{335}{3}-145|\] \[|\frac{10}{3}|+|\frac{335-435}{3}|\] \[|\frac{10}{3}|+|\frac{-100}{3}|=\frac{10}{3}+\frac{100}{3}=\frac{110}{3}\]

myininaya (myininaya):

you do know that if x<0 then |x|=-x and when x>=0 then |x|=x that is why they put negatives in front of the negative parts and positives and front of the positive it is the definition of absolute value

OpenStudy (anonymous):

oh ok so displacement is -30 and distance is 36.6

myininaya (myininaya):

i would leave it as 110/3 instead of 36.7

myininaya (myininaya):

unless you are asked to round to the neartest 10th

myininaya (myininaya):

then it would be 36.7

myininaya (myininaya):

also do you understand the arithmetic now?

OpenStudy (anonymous):

yea thanks and displacement part is -30 right ?

myininaya (myininaya):

that is what we got

OpenStudy (anonymous):

cool thanks for help

myininaya (myininaya):

np